OpenStudy (anonymous):

A student determined the heat of fusion of water to be 88 calories per gram if the accepted value is 80 calories per gram, what is the students percent error?

OpenStudy (anonymous):

it it 11% ???

OpenStudy (xishem):

Percent error is defined as...\[\delta = \frac{\left| v-v_{measured} \right|}{\left| v \right|}*100 \]Where... delta is the percent error v is the actual (or accepted) value v_measured is the value that was measured \[\delta = \frac{\left| 88 - 80 \right|}{\left| 80 \right|}*100=\frac{|8|}{|80|}*100=0.1*100=10\%\]

OpenStudy (anonymous):

Oh Yeah i remember learning that in the beginning of the year! Thank You so much Again! ^.^